# Evaluating the World Model Implicit in a Generative Model

### v2

@@ Abstract
Recent work suggests that large language models may implicitly learn world models. How should we assess this possibility? We formalize this question for the case where the underlying reality is governed by a deterministic finite automaton. This includes problems as diverse as simple logical reasoning, geographic navigation, game-playing, and chemistry. We propose new evaluation metrics for world model recovery inspired by the classic Myhill-Nerode theorem from language theory. We illustrate their utility in three domains: game playing, logic puzzles, and navigation. In all domains, the generative models we consider do well on existing diagnostics for assessing world models, but our evaluation metrics reveal their world models to be far less coherent than they appear. Such incoherence creates fragility: using a generative model to solve related but subtly different tasks can lead it to fail badly. Building generative models that meaningfully capture the underlying logic of the domains they model would be immensely valuable; our results suggest new ways to assess how close a given model is to that goal.
@@ 1 Introduction
To illustrate these ideas, we first take the stylized mapping example literally. We construct a turn-by-turn sequence dataset of taxi rides in New York City. We then assess to what extent transformers (trained on up to 4.7B tokens) successfully recover the true street map of Manhattan. By the usual metrics, the transformers do very well: their predicted next-direction is a valid turn nearly 100% of the time and their state representations even appear to encode the current location of the ride. Our evaluation methods reveal they are very far from recovering the true street map of New York City. As a visualization, we use graph reconstruction techniques to recover each model’s implicit street map of New York City. The resulting map bears little resemblance to the actual streets of Manhattan, containing streets with impossible physical orientations and flyovers above other streets (see Figure 3). Because these transformers fail to recover the true street map of New York City, they are fragile for downstream tasks. While they sometimes have amazing route planning abilities, their performance breaks down when detours are introduced.
@@ 2 Framework
$L^{m}(s)=\{a_{1}a_{2}..a_{k}:(\forall j<k):m(a_{j+1}\leavevmode\nobreak\ |\leavevmode\nobreak\ sa_{1}..a_{j})>0\}.$
@@ 2.1 Recovering world models
Proposition 2.3 (proof given in Appendix A) suggests a way to evaluate whether a generative model recovers the true DFA: assess the validity of next-token predictions. Existing world model diagnostics are motivated by this intuition; for example, Toshniwal et al. [32] and Li et al. [17] assess world model recovery by measuring the percent of top model predictions that are valid in the underlying DFA.
@@ 2.2 Next-token prediction is a fragile metric for recovering structure
Example: Cumulative Connect-4.  Consider a vertical grid with $n$ rows and 7 columns. Two players take turns dropping a disk in a column, and they can choose any column that contains less than $n$ disks. When a disk is dropped in a column, it occupies the bottom-most position that isn’t occupied by another disk, and it remains in that position for the full game. The game continues until the entire board is filled, for $7n$ moves, regardless of whether a player has achieved four in a row. Games are represented as sequences of moves, where each sequence has $7n$ tokens and each token is an integer between 1 and 7 indicating the column the disk is placed in. Here, $\Sigma=\{1,..,7\}$ denotes the columns and the state corresponds to the count in each column. A column is a valid move if that column is not already filled.
[…]
@@ 2.3 The Myhill-Nerode interior and boundary
$\text{MNI}^{W}(q_{1},q_{2})=\{s\in\Sigma^{*}\text{ s.t. }s\in L^{W}(q_{1})\cap L^{W}(q_{2})\}.$
@@
$\text{MNB}^{W}(q_{1},q_{2})=\{s=a_{1}a_{2}..a_{k}\leavevmode\nobreak\ |\leavevmode\nobreak\ s\in L^{W}(q_{1})\setminus L^{W}(q_{2})\text{ and }\forall j<k:a_{1}..a_{j}\in\text{MNI}^{W}(q_{1},q_{2})\}.$
@@
Figure 2: A visual depiction of our two evaluation metrics. A compression error is a model failing to recognize that two sequences that result in the same state should accept the same suffixes. A distinction error is a model failing to find the right distinguishing suffixes for two sequences that lead to different states.
@@ 2.4 Compression and distinction metrics for evaluating world models
$\text{MNB}^{m}(s_{1},s_{2})=\{x=x_{1}..x_{k}|x\in m(s_{1})\setminus m(s_{2})\text{ and }\forall j<k:x_{1}..x_{j}\in m(s_{1})\cap m(s_{2})\}.$ (1)
This is the set of minimal sequences that are accepted by the model conditioned on $s_{1}$ but not $s_{2}$. Since we now focus on the generative model rather than the DFA, the definition refers to pairs of sequences rather than to pairs of states.
@@
$\frac{|\text{MNB}^{W}(q_{1},q_{2})\cap(m(s_{1})\setminus m(s_{2}))|}{|\text{MNB}^{W}(q_{1},q_{2})|},$ (2)
@@
Both our sequence compression and distinction metrics are depicted in Figure 2. Although we have defined a generative model as accepting all sequences it assigns positive mass to, in practice sequence models are regularized to assign all sequences nonzero mass. Our evaluation metrics therefore depend on a parameter $\epsilon\geq 0$, which defines an acceptance threshold for the model. In practice, we explore sensitivity to different values of $\epsilon$. We present parameter ablations and other implementation details in more depth in Section 3 and Appendix D.
@@ 3.1 Data and models
A traversal is a sequence of nodes where an edge exists between each consecutive node in the sequence. To study how the construction of traversals affects the resulting generative model, we consider three different approaches. Shortest paths constructs traversals by finding the shortest path between two nodes. Since these may not be reflective of real-world traversals due to traffic conditions, noisy shortest paths constructs multiple shortest paths by perturbing the magnitude of each edge weight in the underlying graph. Finally, random walks samples random traversals rather than approximating shortest paths. See Appendix E for details.
@@
We randomly split data into train and test splits, ensuring no origin-destination pair is in both train and test sets. We include all sequences containing less than 100 directions. Our training sets consist of 2.9M sequences (120M tokens) for shortest paths; 31M sequences (1.7B tokens) for noisy shortest paths; and 91M sequences (4.7B tokens) for random walks.
We train two types of transformers [34] from scratch using next-token prediction for each dataset: a 117M parameter model consisting of 12 layers, 768 hidden dimensions, and 12 heads; and a 1.6B parameter model consisting of 48 layers, 1600 hidden dimensions, and 25 heads. We follow the architecture of GPT-2 for each model [25]. We train models on 8 A100 GPUs. For each dataset, we analyze the model with the best held-out performance: the 117M parameter model for shortest paths, and the 1.6B parameter for noisy shortest paths and random walks.
@@ 3.2 Evaluating world models
To assess their capabilities, we first assess whether the trained models can recover the shortest paths between unseen (origin, destination) pairs. We prompt each model with (origin, destination) pairs from the test set and use greedy decoding to generate a set of directions.
All models consistently generate valid traversals — between 96 and 99%. Impressively, 97% of the sequences generated by the shortest paths model are the true shortest path, and 94% of the sequences generated by the model trained on noisy shortest paths find a shortest path for one of the noisy graphs used to generate data. Figure 5 provides an example of a shortest path traversal recovered by the shortest path models. Appendix F provides more analysis about traversal capabilities.
To assess whether these capabilities correspond to coherent implicit world models, we first consider two existing diagnostics [32, 17]. The next-token test assesses whether a model, when conditioned on each subsequence in the test set, predicts a legal turn for its top-1 predicted next-token. In our example, a directional move is legal if a street in the direction exists at the current intersection. Predicting the end token is only legal if the traversal implied by the sequence is at the listed destination. Meanwhile, the current-state probe trains a probe [10] from a transformer’s representation to predict the current intersection implied by the directions so far. We train a linear probe on a transformer’s last layer representation and evaluate on held-out data.
To implement the sequence compression metric, we randomly sample states (i.e., [intersection, destination] pairs) and two distinct traversals (i.e. prefixes) that arrive at each state. We then assess whether a model correctly admits the same suffixes for each prefix. We average over pairs of prefixes to report a score for each state and average over states to report a final score. To implement the sequence distinction metrics, we sample pairs of distinct states and traversals (i.e. prefixes) that arrive at each state, comparing the model’s approximate Myhill-Nerode boundary to the true one. We average over pairs of prefixes to report a score for each pair of states, and average over 1000 randomly sampled state pairs to report a final scores. Both metrics depend on a threshold parameter $\epsilon$: a prefix is only sampled or accepted if the model’s assigned probability for each token is above $\epsilon$. Here, we consider $\epsilon=0.01$ for all models and metrics. We describe implementation details and provide parameter ablations in Appendix D.
@@
Untrained transformer 0.03 (0.00) 0.10 (0.00) 0.00 (0.00) 0.00 (0.00) 0.01 (0.00)
Shortest paths 1.00 (0.00) 0.91 (0.00) 0.19 (0.01) 0.36 (0.01) 0.26 (0.01)
Noisy shortest paths 1.00 (0.00) 0.90 (0.00) 0.07 (0.01) 0.36 (0.01) 0.25 (0.01)
Random walks 1.00 (0.00) 0.99 (0.00) 0.68 (0.02) 0.99 (0.00) 1.00 (0.00)
@@
Table 1: Sequence compression and distinction metrics for world models compared to existing metrics (standard errors in parentheses). Models that do well on existing metrics perform poorly on ours.
@@
Our evaluation metrics, however, reveal that these existing diagnostics are incomplete. All trained transformers perform poorly on sequence compression, frequently failing to recognize that two prefixes leading to the same state should admit the same continuations. Even the transformer trained on random walks, which sees many distinct types of traversals during training, fails to compress prefixes for more than a third of states. Appendix F gives a qualitative analysis of states that models fail to compress. For the sequence distinction metrics, the transformers trained on shortest paths or noisy shortest paths perform poorly. In contrast, the transformer trained on random walks performs well on the sequence distinction metric. Both metrics are therefore valuable for evaluating world models; a model can perform well on one metric and poorly on the other. Here, a model that distinguishes separate states at a high rate fails to recognize that two prefixes that lead to the same state should have the same valid continuations.
@@ 3.3 Reconstructing implicit maps
Figure 3 shows three reconstructed maps using sequences generated by the transformer trained on random walks. The sequences underlying each map are generated by randomly sampling 6400 (origin, destination) pairs and then sampling the model’s traversal for each pair. On the left is the reconstructed map on only sequences which are valid under the true world model. On the right is the reconstructed map using the transformer’s sequences. The transformer’s underlying world model is incoherent; it recovers streets whose orientations are physically impossible (e.g. labeled NW but facing east) and require flyovers above other streets.
To show that this map is not the product of a model that has the right world model but makes a few transcription errors, we artificially corrupt sequences drawn from the true model. With probability equal to the probability of an error for the random walks transformer, we randomly re-label an edge in a sequence consistent with the world model. The middle panel of Figure 3 shows the reconstructed graph. It is much closer to the true world model than the transformer (which makes errors at the same rate) is. While these results are for random walks and one setting of graph reconstruction, Appendix B show maps for the other models and considers different settings on graph reconstruction. All settings recover incoherent underlying maps.
@@ 3.4 Implication of failing to recover the world model: detour fragility
Shortest paths 0.99 (0.01) 0.67 (0.05) 0.09 (0.03) 0.00 (0.00) 0.00 (0.00)
Random Noisy shortest paths 0.96 (0.02) 0.74 (0.04) 0.04 (0.02) 0.00 (0.00) 0.00 (0.00)
detours Random walks 0.99 (0.01) 0.99 (0.01) 1.00 (0.00) 0.94 (0.02) 0.80 (0.04)
@@
Shortest paths 0.99 (0.01) 0.63 (0.05) 0.07 (0.03) 0.00 (0.00) 0.00 (0.00)
Adversarial Noisy shortest paths 0.96 (0.02) 0.63 (0.05) 0.07 (0.03) 0.00 (0.00) 0.00 (0.00)
detours Random walks 0.99 (0.01) 0.99 (0.01) 1.00 (0.00) 0.89 (0.03) 0.66 (0.05)
@@ 4 Other Applications: Othello and Logic Puzzles
We apply our metrics to the two Othello sequence models considered by Li et al. [17]: one trained on real games from Othello championship tournaments and another trained on synthetic games. Table 6 in Appendix F shows the result of the metrics in both settings. The model trained on real games performs poorly on both compression and distinction metrics, failing to group together most pairs of game openings that lead to the same board. In contrast, the model trained on synthetic games performs well on both metrics. This distinction is not captured by the existing metrics, which show both models performing similarly. Similar to the navigation setting, we again find that models trained on random/synthetic data recover more world structure than those trained on real-world data.
@@
We analyze Llama 2 (70B) [33], Llama-3 (8B and 70B), Mixtral (8x22B Instruct) [12], Qwen 1.5 Chat (72B and 110B) [2], GPT-3.5 turbo, and GPT-4. We consider $n$=3 individuals here; in Appendix F we show similar results for $n$=5. We first assess whether the LLMs solve the logic puzzle task when the seating arrangement is fully specified by the statements. Figure 4 shows that most LLMs perform well at this task; GPT-4 is accurate on all examples. We then apply our metrics, assessing if each LLM compresses correctly (whether two sets of statements that lead to the same state lead to the same assessments) and has high recall for distinction (we do not compute precision for distinction because it is too expensive to approximate each LLM’s Myhill-Nerode boundary). See Appendix D for further discussion. We perform greedy decoding and encourage each model to perform chain-of-thought reasoning [36].

### v3

@@ Abstract
Recent work suggests that large language models may implicitly learn world models. How should we assess this possibility? We formalize this question for the case where the underlying reality is governed by a deterministic finite automaton. This includes problems as diverse as simple logical reasoning, geographic navigation, game-playing, and chemistry. We propose new evaluation metrics for world model recovery inspired by the classic Myhill-Nerode theorem from language theory. We illustrate their utility in three domains: game playing, logic puzzles, and navigation. In all domains, the generative models we consider do well on existing diagnostics for assessing world models, but our evaluation metrics reveal their world models to be far less coherent than they appear. Such incoherence creates fragility: using a generative model to solve related but subtly different tasks can lead to failures. Building generative models that meaningfully capture the underlying logic of the domains they model would be immensely valuable; our results suggest new ways to assess how close a given model is to that goal.
@@ 1 Introduction
To illustrate these ideas, we first take the stylized mapping example literally. We construct a turn-by-turn sequence dataset of taxi rides in New York City. We then assess to what extent transformers successfully recover the true street map of Manhattan. By the usual metrics, the transformers do very well: their predicted next-direction is a valid turn nearly 100% of the time and their state representations even appear to encode the current location of the ride. Our evaluation methods reveal they are very far from recovering the true street map of New York City. As a visualization, we use graph reconstruction techniques to recover each model’s implicit street map of New York City. The resulting map bears little resemblance to the actual streets of Manhattan, containing streets with impossible physical orientations and flyovers above other streets (see Figure 3). Because these transformers fail to recover the true street map of New York City, they are fragile for downstream tasks. While they sometimes have amazing route planning abilities, their performance breaks down when detours are introduced.
@@ 2 Framework
$L^{m}(s)=\{a_{1}a_{2}...a_{k}:\forall j<k,\ m(a_{j+1}\leavevmode\nobreak\ |\leavevmode\nobreak\ sa_{1}...a_{j})>0\}.$
@@ 2.1 Recovering world models
Proposition 2.3 (proof given in Appendix A) suggests a way to evaluate whether a generative model recovers the true DFA: assess the validity of next-token predictions. Existing world model diagnostics are motivated by this intuition; for example, one way that Toshniwal et al. [36] and Li et al. [20] assess world model recovery is by measuring the percent of top next-token predictions that are valid.
@@ 2.2 Next-token prediction is a fragile metric for recovering structure
Example: Cumulative Connect-4.  Consider a vertical grid with $n$ rows and 7 columns. Two players take turns dropping a disk in a column, and they can choose any column that contains less than $n$ disks. When a disk is dropped in a column, it occupies the bottom-most position that isn’t occupied by another disk, and it remains in that position for the full game. The game continues until the entire board is filled, for $7n$ moves, regardless of whether a player has achieved four in a row. Games are represented as sequences of moves, where each sequence has $7n$ tokens and each token is an integer between 1 and 7 indicating the column the disk is placed in. Here, $\Sigma=\{1,\dots,7\}$ denotes the columns and the state corresponds to the count in each column. A column is a valid move if that column is not already filled.
Consider a generative model that outputs $\{1,\dots,7\}$ with uniform probability given any sequence, i.e. $m(a\leavevmode\nobreak\ |\leavevmode\nobreak\ s)=m(a^{\prime}\leavevmode\nobreak\ |\leavevmode\nobreak\ s^{\prime})=1/7$ for all $a,a^{\prime}\in\Sigma$ and $s,s^{\prime}\in\Sigma^{*}$. This model clearly encodes no information about the board. However, for any board where there are no columns filled, this model provides a valid next move (e.g., the right panel of Figure 1), and so it will be a near-perfect next-token predictor when $n$ is large. For example, when $n=1000$, it predicts a valid next move for more than 99% of all states. Metrics based on next-token prediction will imply this algorithm is close to recovering a world model.
@@ 2.3 The Myhill-Nerode interior and boundary
$\text{MNI}^{W}(q_{1},q_{2})=\{s\in\Sigma^{*}\leavevmode\nobreak\ |\leavevmode\nobreak\ s\in L^{W}(q_{1})\cap L^{W}(q_{2})\}.$
@@
$\text{MNB}^{W}(q_{1},q_{2})=\{s=a_{1}a_{2}...a_{k}\leavevmode\nobreak\ |\leavevmode\nobreak\ s\in L^{W}(q_{1})\setminus L^{W}(q_{2})\text{ and }\forall j<k:a_{1}...a_{j}\in\text{MNI}^{W}(q_{1},q_{2})\}.$
@@
Figure 2: A visual depiction of our two evaluation metrics. A compression error is a model failing to recognize that two sequences that result in the same state should accept the same suffixes. A distinction error is a model failing to find the right distinguishing suffixes for two sequences that lead to different states. Our metrics measure errors at the boundary, which are visually depicted above.
@@ 2.4 Compression and distinction metrics for evaluating world models
$\text{MNB}^{m}(s_{1},s_{2})=\{x=x_{1}...x_{k}\leavevmode\nobreak\ |\leavevmode\nobreak\ x\in L^{m}(s_{1})\setminus L^{m}(s_{2})\text{ and }\forall j<k:x_{1}...x_{j}\in L^{m}(s_{1})\cap L^{m}(s_{2})\}.$ (1)
This is the set of minimal suffixes that are accepted by the model conditioned on $s_{1}$ but not $s_{2}$. Since we now focus on the generative model rather than the DFA, the definition refers to pairs of sequences rather than to pairs of states.
@@
$\frac{|\text{MNB}^{W}(q_{1},q_{2})\cap(L^{m}(s_{1})\setminus L^{m}(s_{2}))|}{|\text{MNB}^{W}(q_{1},q_{2})|},$ (2)
@@
Both metrics are depicted in Figure 2. Although we have defined a generative model as accepting all sequences it assigns positive probability to, in practice sequence models are regularized to assign all sequences nonzero probability. Our evaluation metrics therefore depend on an acceptance threshold parameter $\epsilon>0$. In practice, we explore sensitivity to different values of $\epsilon$ and other acceptance mechanisms. We present ablations and other details in more depth in Section 3 and Appendix E.
@@ 3.1 Data and models
A traversal is a sequence of nodes where an edge exists between each consecutive node in the sequence. To study how the construction of traversals affects the resulting generative model, we consider three different approaches. Shortest paths constructs traversals by finding the shortest path between two nodes. Since these may not be reflective of real-world traversals due to traffic conditions, noisy shortest paths constructs multiple shortest paths by perturbing the magnitude of each edge weight in the underlying graph. Finally, random walks samples random traversals instead of approximating shortest paths. See Appendix F for details.
@@
We randomly split data into train and test splits, ensuring no origin-destination pair is in both train and test sets. We include all sequences containing less than 100 directions. Our training sets consist of 2.9M sequences (120M tokens) for shortest paths; 31M sequences (1.7B tokens) for noisy shortest paths; and 91M sequences (4.7B tokens) for random walks. We train two types of transformers [38] from scratch using next-token prediction for each dataset: an 89.3M parameter model consisting of 12 layers, 768 hidden dimensions, and 12 heads; and a 1.5B parameter model consisting of 48 layers, 1600 hidden dimensions, and 25 heads. We follow the architecture of GPT-2 for each model [29]. We train models on 8 A100 GPUs. For each dataset, we analyze the model with the best held-out performance: the 89.3M parameter model for shortest paths, and the 1.5B parameter for noisy shortest paths and random walks.
@@ 3.2 Evaluating world models
To assess their capabilities, we first assess whether the trained models can recover the shortest paths between unseen (origin, destination) pairs. We prompt each model with (origin, destination) pairs from the test set and use greedy decoding to generate a set of directions. All models consistently generate valid traversals — between 96% and 99%. Impressively, 97% of the sequences generated by the shortest paths model are the true shortest path, and 94% of the sequences generated by the model trained on noisy shortest paths find a shortest path for one of the noisy graphs used to generate data. Figure 5 provides an example of a shortest path traversal.
To assess whether these capabilities correspond to coherent implicit world models, we first consider two existing diagnostics [36, 20]. The next-token test assesses whether a model, when conditioned on each subsequence in the test set, predicts a legal turn for its top-1 predicted next-token. In our example, a directional move is legal if a street in the direction exists at the current intersection. Predicting the end token is only legal if the traversal implied by the sequence is at the listed destination. Meanwhile, the current-state probe trains a probe [11] from a transformer’s representation to predict the current intersection implied by the directions so far. We train a linear probe on a transformer’s last layer representation.
To implement the sequence compression metric, we randomly sample states (i.e., [intersection, destination] pairs) and two distinct traversals (i.e. prefixes) that arrive at each state. We then assess whether a model correctly admits the same suffixes for each prefix. We average over pairs of prefixes to report a score for each state and average over states to report a final score. To implement the sequence distinction metrics, we sample pairs of distinct states and traversals (i.e. prefixes) that arrive at each state, comparing the model’s approximate Myhill-Nerode boundary to the true one. We average over pairs of prefixes to report a score for each pair of states, and average over 1000 randomly sampled state pairs to report a final scores. Both metrics depend on a threshold parameter $\epsilon$: a prefix is only sampled or accepted if the model’s assigned probability for each token is above $\epsilon$. Here, we consider $\epsilon=0.01$ for all models and metrics. We describe implementation details, provide parameter ablations, and consider other acceptance rules (e.g. top-p and top-k) in Appendix E.
@@
Untrained transformer 0.03 (0.00) 0.10 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Shortest paths 1.00 (0.00) 0.91 (0.00) 0.10 (0.01) 0.35 (0.02) 0.20 (0.01)
Noisy shortest paths 1.00 (0.00) 0.92 (0.00) 0.05 (0.01) 0.37 (0.02) 0.24 (0.01)
Random walks 1.00 (0.00) 0.99 (0.00) 0.50 (0.02) 0.99 (0.00) 1.00 (0.00)
@@
Table 1: Sequence compression and distinction metrics for world models compared to existing metrics (standard errors in parentheses). Models that do well on existing metrics can perform poorly on ours.
@@
Our evaluation metrics, however, reveal that these existing diagnostics are incomplete. All trained transformers perform poorly on sequence compression, frequently failing to recognize that two prefixes leading to the same state should admit the same continuations. Even the transformer trained on random walks, which sees many distinct types of traversals during training, fails to compress prefixes for half the states. For the sequence distinction metrics, the transformers trained on shortest paths or noisy shortest paths perform poorly. In contrast, the transformer trained on random walks performs well on the sequence distinction metric. Both metrics are therefore valuable for evaluating world models; a model can perform well on one metric and poorly on the other. Here, a model that distinguishes separate states at a high rate fails to recognize that two prefixes that lead to the same state should have the same valid continuations.
@@ 3.3 Reconstructing implicit maps
Figure 3 shows three reconstructed maps using sequences generated by the transformer trained on random walks. The sequences underlying each map are generated by randomly sampling 6400 (origin, destination) pairs and then sampling the model’s traversal for each pair (Appendix G shows similar results for when the distribution of origin/destination pairs follows the sampling distribution used to train each model). On the left is the reconstructed map on only sequences which are valid under the true world model. On the right is the reconstructed map using the transformer’s sequences. The transformer’s underlying world model is incoherent; it recovers streets whose orientations are physically impossible (e.g. labeled NW but facing east) and require flyovers above other streets.
To show that this map is not the product of a model that has the right world model but makes a few transcription errors, we artificially corrupt sequences drawn from the true model. With probability equal to the probability of an error for the random walks transformer, we randomly re-label an edge in a sequence consistent with the world model. The middle panel of Figure 3 shows the reconstructed graph. It is much closer to the true world model than the transformer (which makes errors at the same rate). While these results are for random walks and one setting of graph reconstruction, Appendix G shows maps for the other models and different reconstruction settings. All settings recover incoherent underlying maps.
@@ 3.4 Implication of failing to recover the world model: detour fragility
Shortest paths 0.99 (0.01) 0.69 (0.05) 0.08 (0.03) 0.00 (0.00) 0.00 (0.00)
Random Noisy shortest paths 0.96 (0.02) 0.52 (0.05) 0.03 (0.02) 0.00 (0.00) 0.00 (0.00)
detours Random walks 0.99 (0.01) 0.99 (0.01) 1.00 (0.00) 0.97 (0.02) 0.74 (0.04)
@@
Shortest paths 0.99 (0.01) 0.66 (0.05) 0.06 (0.02) 0.00 (0.00) 0.00 (0.00)
Adversarial Noisy shortest paths 0.96 (0.02) 0.64 (0.05) 0.04 (0.02) 0.00 (0.00) 0.00 (0.00)
detours Random walks 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.93 (0.03) 0.51 (0.05)
@@ 4 Other Applications: Othello and Logic Puzzles
We apply our metrics to the two Othello sequence models considered by Li et al. [20]: one trained on real games from Othello championship tournaments and another trained on synthetic games. Table 3 in Appendix D reports the metrics in both settings. The model trained on real games performs poorly on both compression and distinction metrics, failing to group together most pairs of game openings that lead to the same board. In contrast, the model trained on synthetic games performs well on both metrics. This discernment is not captured by the existing metrics, which show both models performing similarly. We validate this discernment by performing a “detours” exercise for Othello in Table 4 in Appendix D; while the model trained on synthetic data produces near-perfect games regardless of detours, the model trained on championship data fails immediately. Similar to the navigation setting, we again find that models trained on random/synthetic data recover more structure than those trained on real-world data.
@@
We analyze Llama 2 (70B) [37], Llama-3 (8B and 70B), Mixtral (8x22B Instruct) [15], Qwen 1.5 Chat (72B and 110B) [2], GPT-3.5 turbo, and GPT-4. We consider $n$=3 individuals. We first assess whether the LLMs solve the logic puzzle task when the seating arrangement is fully specified by the statements. Figure 4 shows that most LLMs perform well at this task; GPT-4 is accurate on all examples. We then apply our metrics, assessing if each LLM compresses correctly (whether two sets of statements that lead to the same state lead to the same assessments) and has high recall for distinction (we do not compute precision for distinction because it is too expensive to approximate each LLM’s Myhill-Nerode boundary). See Appendix E for further discussion.